# PsycoLLM: Enhancing LLM for Psychological Understanding and Evaluation

### v2

@@ Abstract
Mental health has attracted substantial attention in recent years and LLM can be an effective technology for alleviating this problem owing to its capability in text understanding and dialogue. However, existing research in this domain often suffers from limitations, such as training on datasets lacking crucial prior knowledge and evidence, and the absence of comprehensive evaluation methods. In this paper, we propose a specialized psychological large language model (LLM), named PsycoLLM, trained on a proposed high-quality psychological dataset, including single-turn QA, multi-turn dialogues and knowledge-based QA. Specifically, we construct multi-turn dialogues through a three-step pipeline comprising generation, evidence judgment, and refinement. We augment this process with real-world psychological case backgrounds extracted from online platforms, enhancing the relevance and applicability of the generated data. Additionally, to compare the performance of PsycoLLM with other LLMs, we develop a comprehensive psychological benchmark based on authoritative psychological counseling examinations in China, which includes assessments of professional ethics, theoretical proficiency, and case analysis. The experimental results on the benchmark illustrates the effectiveness of PsycoLLM, which demonstrates superior performance compared to other LLMs.
@@ I Introduction
Large language models (LLM) have attracted tremendous attention and experienced promising advancements, such as ChatGPT [^1] https://chat.openai.com/, KimiChat [^2] https://kimi.moonshot.cn/, PaLM [1] , BaiChuan [2] and LLama [3]. These models have demonstrated significant advancements across a variety of tasks, including question answering (QA), summarization [4], and information extraction [5]. Moreover, large language models (LLMs) have been effectively applied to diverse domains, such as the medical and legal domains [6, 7, 8]. In recent years, the escalating pressures from personal, and occupational demands have emerged as significant challenges to the mental health status of individuals. Since LLMs are trained on massive data and contain huge number of parameters, many experiments have shown that these models have the strong ability to understand the natural language and provide resaonable feedback. Hence, leveraging LLM to aid individuals in addressing psychological issues may represent an efficacious strategy for mitigating the impact of adverse mental health conditions and making it easier for seekers to obtain timely help.
Many studies have been proposed in this area. For example, The work [9] combined pre-trained LLMs with real-world professional QA data, with the purpose of providing online consultation services. [10] and [11] constructed a multi-turn empathetic conversation dataset and used a base LLM for fine-tuning. Further, EmoLLM [12] follow a similar way and fine-tune the open base model on the mental health dataset. Although these existing efforts have contributed significantly to the advancement of the field, they still exhibit certain limitations. First, most existing psychological fine-tuning datasets are generated by even larger language model such as ChatGPT. However, the generation process lacks adequate prior knowledge and evidence, potentially leading to the creation of datasets that deviate from real-world psychological communication scenarios. Second, owing to the lack of a comprehensive psychological benchmark, the evaluation of most studies in this area often relies on expert assessment based on subjective judgment or overly depends on other large models, which may not fully encompass the subtle performance. For instance, PsyBench [13] utilized questions generated by GPT-4 to evaluate model performance.
@@
The results reveal several insightful observations: I) Some LLMs have achieved an accuracy rate surpassing 60% in MCQs, showing the potential of using LLM to provide psychological assistance, with the proposed PsycoLLM demonstrating superior performance among these models. II) LLMs display greater proficiency in professional ethics compared to their mastery of psychological theory.
@@
  • We propose a psychological LLM (PsycoLLM), fine-tuned on the high-quality psychological dataset.
@@ II Related Work
Figure 1: Overview of dataset preparation.
@@ II-A Large Language Model
Owing to highly parallelizable architecture of self-attention mechanisms, Transformer-based models have been popular in recent years. BERT [14] was a first pre-trained Transformer-based language model, which trained a bidirectional language encoder with specially designed masked language modelling. GPT [15] trained a Transformer-based decoder by generative pre-training. Subsequently, GPT-2 [16] was introduced as an expanded version of GPT, featuring a similar architecture but with a significantly larger scale, encompassing up to 1.5 billion parameters. BART [17] used a sequence-to-sequence architecture and combined the strengths of BERT and GPT, which employed a bidirectional encoder similar to BERT and an autoregressive decoder similar to GPT. These models typically followed a pre-training and fine-tuning paradigm, which need to fine-tune these model for different downstream tasks. Owing the findings that the larger models typically yield better performance, LLM become a prominent topic of research and have achieved significant success. For example, GPT-3 [18] with 175B parameters surpassed its predecessor, GPT-2, and performed well in few-shot learning. Recently, ChatGPT based on GPT-3.5, trained with a sizable amount of internet-sourced text data, has drawn sustainable attention which is exceptional at producing human-like responses. Building upon this success, OpenAI introduced GPT-4 [19], which further enhanced the capabilities of its predecessors. In contrast to ChatGPT and GPT-4 were proprietary models, LLaMA [3] was a open-resource language model, which attracted widespread attention from both academia and industry due to its accessible nature and has achieved noteworthy performance. Further, Touvron et al [20] proposed an updated version of LLamA, called LLaMA2, which was trained on a larger corpus and outperformed LLaMA1 in many benchmarks. Additionally, many other LLMs based on LLaMA architecture also have been proposed, including Alpaca [21] and Mistral-7B [22], where the former first employed instruct-following to train the LLaMA and the latter is a 7B-parameter and outperformed LLamA-2-13B in various evaluations. Baichuan [2] introduced a series of large-scale multilingual language models containing 7 billion and 13 billion parameters, trained from scratch on a corpus of 2.6 trillion tokens. ChatGLM [23] was an evolving family of large language models, which included ChatGLM [24, 25], ChatGLM2 and ChatGLM3, repectively. Besides, the Yi series models [26] were also trained from scratch on 3T multilingual corpus and notable for their ability to extend context length up to 200K by continuing pretraining. Furthermore, Qwen [27] represented another open-resource LLMs and obtained superior performance across a multitude of downstream tasks, especially in Chinese, which was also the main reason we choose it as our backbone.
Figure 2: The topic distribution.
@@ II-B Benchmark for Mental Health Support
Many studies have been devoted to evaluating the performance of the model across various domains. For example, in the general domain, [28] proposed a Chinese evaluation benchmark to evaluate advanced knowledge and reasoning abilities of Chinese-based models. Zhong at al [29] designed a human-centric benchmark for assessing the performance of foundation models via 20 diverse tasks across a wide variety of subjects. The MMLU benchmark [30] provided a comprehensive evaluation benchmark encompassing various tasks to evaluate the performance of a text model in multi-task contexts. In the medical domain, [31] proposed a localized medical benchmark across various clinical medical professions to evaluate the existing LLMs for medicine. In the legal domain, [32] constructed a Chinese legal benchmark based on the logic of legal practice, categorizing the legal capabilities of language models into different levels. In the psychological domain, several studies have introduced benchmarks to assess model performance in mental health contexts. Jin et al [33] introduced a meticulously crafted benchmark to evaluate the model performance in five mental health-related sub-tasks. Zhang et al [13] used GPT-4 to generate question and constructed a psychological benchmark aimed at evaluating model performance in psychology. Furthermore, Qiu et al [34] created a benchmark with fine-grained labels for each dialogue and focused on dialogue safety in mental health support. Despite the advances made with these benchmarks, there are several limitations. Some benchmarks narrow their scope to specific subfields or isolated tasks within mental health, neglecting to evaluate the models’ overall psychological knowledge mastery. Furthermore, benchmarks generated by large models carry the risk of deviating from real-world contexts. Compared to these works, we propose a benchmark that incorporates different levels of official examinations in China, enabling a comprehensive assessment of both theoretical knowledge and practical proficiency.
@@ II-C Mental Application in LLM
The success of LLMs in general domains has sparked interest in their application across various vertical fields. Many LLMs tailored for specific domains have achieved remarkable success [35, 36]. For instance, HuatuoGPT [37], an LLM designed for the medical domain, has demonstrated excellent performance in medical consultations. Based on HuatuoGPT, HuatuoGPT-II [38] introduced a unified domain adaptation protocol to combine continued pre-training and supervised fine-tuning into a single process, which achieved better performance than HuatuoGPT. In the legal domain, Yue et al [39] developed an LLM tailored for building intelligent legal systems with legal reasoning and knowledge retrieval capability. Mental health is another crucial domain that has garnered significant attention in many studies [40, 41, 42, 43, 44, 45, 46]. For example, the work [47] developed an LLM-based expert-guided pipeline to analyze raw interview transcripts and extract insights. Zhu et al [48] investigated the use of LLM to perform inferring users’ underlying goals and fundamental psychological needs, finding that LLM-based approaches can be comparable to human performance. Wu et al [49] introduced an anthropomorphic LLM-based conversational agent designed to offer personalized support and suggest practical actions based on positive psychology and social psychology research.
Several LLMs have also been proposed to aid help-seekers in accessing effective mental health support. Lai et al [9] proposed an AI-based assistive tool that utilized LLMs for QA, aiming to alleviate the demand on mental health professionals. PsyQA [50] introduced a high-quality Chinese dataset of psychological health support in the form of question and answer pair. Based on PsyQA, Qiu et al [11] further used ChatGPT to covert single-turn dialogues to multi-turn dialogues and subsequently fine-tuned ChatGLM2-6B on this enhanced dataset. SoulChat [10] constructed a multi-turn empathetic conversation dataset and fine-tuned ChatGLM with this data. Yang et al [44] formalized the interpretable mental health analysis task and built a multi-task and multi-source interpretable mental health instruction dataset. After that, based on the contructed dataset, they released MentaLLaMA, an open-source instruction-following LLM fine-tuned on LLaMA2. The work [51] fine-tuned the LLM across multiple datasets simultaneously, demonstrating its effectiveness in various mental health prediction tasks. Although these efforts have greatly advanced the field, they rely too heavily on distillations from other models during data generation, leading to the creation of data that may not accurately reflect real-world scenarios. In this paper, we constructed a high-quality psychological dataset, including single-turn, multi-turn dialogue data, and knowledge QA data.
@@ III PsycoLLM
In this section, we present the details of dataset including, single-turn QA, multi-turn dialogue and knowledge-based QA, respectively. The overview of dataset are shown in Fig. 1. Besides, we give the training process of PsycoLLM.
@@ III-A Single-Turn QA Construction
There are several publicly accessible websites committed to establishing a psychology platform and offering online solutions for individuals seeking psychological assistance, such as Yixinli, Zhihu and so on. The data from these platforms can be regarded as real-world inquiries and the solutions provided by professionals in response to these inquiries. On the original website, users typically provide descriptions along with their questions, including background information. Subsequently, professional or experienced individuals offer detailed solutions or advice in response to these questions. We collect over 267,000 pairs of data from websites. Furthermore, several important data cleaning procedures are implemented to enhance the quality and relevance of the dataset, including:
@@
After these data cleaning procedures, we obtain over 155k pairs for single-turn QA. We report detailed statistics of the processed data drawn from different aspects. We divide the processed data into 9 major topics and 127 subtopics, as illustrated by the distribution of major topics depicted in Fig. 2. It is observed that the number of reports related to emotional issues and regulation, as well as interpersonal relationships and socializing, is higher than those in other categories, each comprising more than 20% of the total. Following these topics, family and marriage, and personal growth and development are the next most significant areas, each accounting for more than 10% of the reports. Besides, the visualization of the collected data is depicted through word clouds, as shown in Fig. 3, which provides an intuitive depiction of the frequency and prominence of specific terms within the dataset. The analysis underscores the pervasive connection of the majority of these data to everyday psychological counseling, further emphasizing the relevance and reasonability of the collected data.
@@
You are an expert who has studied countless conversations between patients with mental health problems and mental health doctors…
Please construct a continuous multi round dialogue record between help-seekers and psychologist that is in line with the actual situation.
@@
You are a psychology professor and are proficient in psychology and linguistics. …
From the perspective of a professional psychology professor, please make a strict judgment based on the given four indicators.
1. Empathy, whether the psychologist can truly understand the emotions and needs of the visitor, express sincere care and understanding …
@@ 4. Safety, whether it may have a negative impact on visitors…
# Avg. Turns per Dialogue 5.90
# Avg. Tokens per Turn 136.28
# Avg. Tokens per Question 43.62
# Avg. Tokens per Answer 92.67
@@ III-B Multi-Turn Dialogue Construction
III-C knowledge-based QA
In addition to the daily QA interactions, we also introduce some more abstract psychological knowledge data, such as explanations of psychological terms. We crawl books related to psychology from the web and then use Qwen-72B to extract knowledge-based QA from them. Specifically, we segment books into text spans using a predefined fixed length, identifying the nearest sentence or paragraph as segmentation indicators. These text spans serve as the fundamental units for subsequent QA generation through the utilization of LLMs. First, the LLM generates questions and their corresponding answers. These question-answer pairs are then input into two LLM-based student modules, one utilizing retrieval-augmented generation (RAG) and the other without RAG, to produce two new sets of answers. Subsequently, a teacher module, also based on an LLM, evaluates and selects best answer from those generated by the student modules. Furthermore, to ensure the quality and accuracy of the generated QA pairs, a manual validation process is implemented, wherein human evaluators assess and eliminate low-quality data. In addition, we extract after-school exercises from several books and convert them to QA format, with the corresponding answer analyses. Finally, we obtain 10K knowledge QA data.

### v3

@@ Abstract
Mental health has attracted substantial attention in recent years and LLM can be an effective technology for alleviating this problem owing to its capability in text understanding and dialogue. However, existing research in this domain often suffers from limitations, such as training on datasets lacking crucial prior knowledge and evidence, and the absence of comprehensive evaluation methods. In this paper, we propose a specialized psychological large language model (LLM), named PsycoLLM, trained on a proposed high-quality psychological dataset, including single-turn QA, multi-turn dialogues and knowledge-based QA. Specifically, we construct multi-turn dialogues through a three-step pipeline comprising multi-turn QA generation, evidence judgment, and dialogue refinement. We augment this process with real-world psychological case backgrounds extracted from online platforms, enhancing the relevance and applicability of the generated data. Additionally, to compare the performance of PsycoLLM with other LLMs, we develop a comprehensive psychological benchmark based on authoritative psychological counseling examinations in China, which includes assessments of professional ethics, theoretical proficiency, and case analysis. The experimental results on the benchmark illustrate the effectiveness of PsycoLLM, which demonstrates superior performance compared to other LLMs. [^1] https://github.com/MACLAB-HFUT/PsycoLLM
@@ I Introduction
Large language models (LLM) have attracted tremendous attention and experienced promising advancements, such as ChatGPT [^2] https://chat.openai.com/, KimiChat [^3] https://kimi.moonshot.cn/, PaLM [1] , BaiChuan [2] and LLama [3]. These models have demonstrated significant advancements across a variety of tasks, including question answering (QA), summarization [4], and information extraction [5]. Moreover, LLMs have been effectively applied to diverse domains, such as the medical and legal domains [6, 7, 8]. In recent years, the escalating pressures from personal, and occupational demands have emerged as significant challenges to the mental health status of individuals. Since LLMs are trained on massive data and contain a huge number of parameters, many experiments have shown that these models have the strong ability to understand the natural language and provide reasonable feedback. Hence, thanks to its advanced text understanding capabilities, leveraging LLM to aid individuals in addressing psychological issues may represent an efficacious strategy for mitigating the impact of adverse mental health conditions and making it easier for seekers to obtain timely help.
Many studies have been proposed in this area. For example, the work [9] combined pre-trained LLMs with real-world professional QA data, with the purpose of providing online consultation services. [10] and [11] constructed a multi-turn empathetic conversation dataset and used a base LLM for fine-tuning. Further, EmoLLM [12] followed a similar way and fine-tuned the open base model on the mental health dataset. Although these existing efforts have contributed significantly to the advancement of the field, they still exhibit certain limitations. First, most existing psychological fine-tuning datasets are generated by even larger language models such as ChatGPT. However, the generation process lacks adequate prior knowledge and evidence, potentially leading to the creation of datasets that deviate from real-world psychological communication scenarios. Second, owing to the lack of a comprehensive psychological benchmark, the evaluation of most studies in this area often relies on expert assessment based on subjective judgment or overly depends on other large models, which may not fully encompass the subtle performance. For instance, PsyBench [13] utilized questions generated by GPT-4 to evaluate model performance.
@@
The results reveal several insightful observations: I) Some LLMs have achieved an accuracy rate surpassing 60% in MCQs, indicating their potential to excel in psychological counseling examinations, with the proposed PsycoLLM demonstrating superior performance among these models. II) LLMs display greater proficiency in professional ethics compared to their mastery of psychological theory.
@@
  • We introduce PsycoLLM, a LLM fine-tuned on the high-quality psychological dataset.
@@ II Related Work
Figure 1: Overview of dataset preparation, including single-turn QA, multi-turn dialogue, and knowledge QA generation.
@@ II-A Large Language Model
Owing to highly parallelizable architecture of self-attention mechanisms, Transformer-based models have been popular in recent years. BERT [14] was a first pre-trained Transformer-based language model, which trained a bidirectional language encoder with specially designed masked language modeling. GPT [15] trained a Transformer-based decoder by generative pre-training. Subsequently, GPT-2 [16] was introduced as an expanded version of GPT, featuring a similar architecture but with a significantly larger scale, encompassing up to 1.5 billion parameters. BART [17] used a sequence-to-sequence architecture and combined the strengths of BERT and GPT, which employed a bidirectional encoder similar to BERT and an autoregressive decoder similar to GPT. These models typically followed a pre-training and fine-tuning paradigm, which need to fine-tune these model for different downstream tasks. Owing the findings that the larger models typically yield better performance, LLM become a prominent research topic and have achieved significant success. For example, GPT-3 [18] with 175B parameters surpassed its predecessor, GPT-2, and performed well in few-shot learning.
Recently, ChatGPT based on GPT-3.5, trained with a sizable amount of internet-sourced text data, has drawn sustainable attention which is exceptional at producing human-like responses. Building upon this success, OpenAI introduced GPT-4 [19], which further enhanced the capabilities of its predecessors. In contrast to ChatGPT and GPT-4 were proprietary models, LLaMA [3] was an open-resource language model, which attracted widespread attention from both academia and industry due to its accessible nature and has achieved noteworthy performance. Further, Touvron et al. [20] proposed an updated version of LLaMA, called LLaMA2, which was trained on a larger corpus and outperformed LLaMA1 in many benchmarks. Additionally, many other LLMs based on LLaMA architecture also have been proposed, including Alpaca [21] and Mistral-7B [22], where the former first employed instruct-following to train the LLaMA and the latter is a 7B-parameter and outperformed LLamA-2-13B in various evaluations. Baichuan [2] introduced a series of large-scale multilingual language models containing 7 billion and 13 billion parameters, trained from scratch on a corpus of 2.6 trillion tokens. ChatGLM [23] was an evolving family of large language models, which included ChatGLM [24, 25], ChatGLM2 and ChatGLM3, repectively. Besides, the Yi series models [26] were also trained from scratch on 3 trillion multilingual corpus and notable for their ability to extend context length up to 200K by continuing pretraining. Furthermore, Qwen [27] represented another open-resource LLMs and obtained superior performance across a multitude of downstream tasks, especially in Chinese, which was also the main reason we choose it as our backbone.
Figure 2: The topic distribution. We divide the data into 9 topics and report their percentages.
@@ II-B Benchmark for Mental Health Support
Many studies have been devoted to evaluating the performance of the model across various domains. For example, in the general domain, [28] proposed a Chinese evaluation benchmark to evaluate the advanced knowledge and reasoning abilities of Chinese-based models. Zhong et al. [29] designed a human-centric benchmark for assessing the performance of foundation models via 20 diverse tasks across a wide variety of subjects. The MMLU benchmark [30] provided a comprehensive evaluation benchmark encompassing various tasks to evaluate the performance of a text model in multi-task contexts. In the medical domain, [31] proposed a localized medical benchmark across various clinical medical professions to evaluate the existing LLMs for medicine. In the legal domain, [32] constructed a Chinese legal benchmark based on the logic of legal practice, categorizing the legal capabilities of language models into different levels.
In the psychological domain, several studies have introduced benchmarks to assess model performance in mental health contexts. Jin et al. [33] introduced a meticulously crafted benchmark to evaluate the model performance in five mental health-related sub-tasks. Zhang et al. [13] used GPT-4 to generate question and constructed a psychological benchmark aimed at evaluating model performance in psychology. Furthermore, Qiu et al. [34] created a benchmark with fine-grained labels for each dialogue and focused on dialogue safety in mental health support. Despite the advances made with these benchmarks, there are several limitations. Some benchmarks narrow their scopes to specific subfields or isolated tasks within mental health, neglecting to evaluate the models’ overall psychological knowledge mastery. Furthermore, benchmarks generated by large models carry the risk of deviating from real-world contexts. Compared to these works, we propose a benchmark that incorporates different levels of official examinations in China, enabling a comprehensive assessment of both theoretical knowledge and practical proficiency.
@@ II-C Mental Application in LLM
Many traditional deep learning-based studies have been proposed for several domains [35, 36, 37, 38, 39, 40]. Recently, the success of LLMs in general domains has sparked interest in their application across various vertical fields. Many LLMs tailored for specific domains have achieved remarkable success [41, 42, 43, 44]. For instance, HuatuoGPT [45], an LLM designed for the medical domain, has demonstrated excellent performance in medical consultations. Based on HuatuoGPT, HuatuoGPT-II [46] introduced a unified domain adaptation protocol to combine continued pre-training and supervised fine-tuning into a single process, which achieved better performance than HuatuoGPT. In the legal domain, Yue et al. [47] developed an LLM tailored for building intelligent legal systems with legal reasoning and knowledge retrieval capability. Mental health is another crucial domain that has garnered significant attention in many studies [48, 49, 50, 51, 52, 53, 54]. For example, the work [55] developed an LLM-based expert-guided pipeline to analyze raw interview transcripts and extract insights. Zhu et al. [56] investigated the use of LLM to perform inferring users’ underlying goals and fundamental psychological needs, finding that LLM-based approaches can be comparable to human performance. Wu et al. [57] introduced an anthropomorphic LLM-based conversational agent designed to offer personalized support and suggest practical actions based on positive psychology and social psychology research.
Several LLMs have also been proposed to aid help-seekers in accessing effective mental health support. Lai et al. [9] proposed an AI-based assistive tool that utilized LLMs for QA, aiming to alleviate the demand on mental health professionals. PsyQA [58] introduced a high-quality Chinese dataset of psychological health support in the form of question and answer pairs. Based on PsyQA, Qiu et al. [11] further used ChatGPT to convert single-turn dialogues to multi-turn dialogues and subsequently fine-tuned ChatGLM2-6B on this enhanced dataset. SoulChat [10] constructed a multi-turn empathetic conversation dataset and fine-tuned ChatGLM with this data. Yang et al. [52] formalized the interpretable mental health analysis task and built a multi-task and multi-source interpretable mental health instruction dataset. After that, based on the contructed dataset, they released MentaLLaMA, an open-source instruction-following LLM fine-tuned on LLaMA2. The work [59] fine-tuned the LLM across multiple datasets simultaneously, demonstrating its effectiveness in various mental health prediction tasks. Although these efforts have greatly advanced the field, they rely too heavily on distillations from other models during data generation, leading to the creation of data that may not accurately reflect real-world scenarios. In this paper, we constructed a high-quality psychological dataset, including single-turn, multi-turn dialogue data, and knowledge QA data.
@@ III PsycoLLM
In this section, we present the details of dataset, including single-turn QA, multi-turn dialogue and knowledge-based QA, respectively. The overview of dataset are shown in Fig. 1. Besides, we give the training process of PsycoLLM.
@@ III-A Single-Turn QA Construction
There are several publicly accessible websites committed to establishing a psychology platform and offering online solutions for individuals seeking psychological assistance, such as Yixinli [^4] https://www.xinli001.com/, Zhihu [^5] https://www.zhihu.com/ and so on. The data from these platforms can be regarded as real-world inquiries and the solutions provided by professionals in response to these inquiries. On the original website, users typically provide descriptions along with their questions, including background information. Subsequently, professional or experienced individuals offer detailed solutions or advice in response to these questions. We collect over 267,000 pairs of data from websites. Furthermore, several important data cleaning procedures are implemented to enhance the quality and relevance of the dataset, including:
@@
You are an expert who has studied countless conversations between patients with mental health problems and doctors…
Please construct a continuous multi round dialogue record between help-seekers and psychologist …
@@
You are a psychology professor and are proficient in psychology and linguistics …
From the perspective of a psychology professor, please make a strict judgment based on the given four indicators.
1. Empathy, whether the psychologist can truly understand the emotions and needs of the visitor, express sincere care …
@@ 4. Safety, whether it may have a negative impact on visitors …
After these data cleaning procedures, we obtain over 155k pairs for single-turn QA. We report detailed statistics of the processed data drawn from different aspects. We divide the processed data into 9 major topics, as illustrated by the distribution of major topics depicted in Fig. 2. It is observed that the number of reports related to emotional issues and regulation, as well as interpersonal relationships and socializing, is higher than those in other categories, each comprising more than 20% of the total. Following these topics, family and marriage, and personal growth and development are the next most significant areas, each accounting for more than 10% of the reports. Besides, the visualization of the collected data is depicted through word clouds, as shown in Fig. 3, which provides an intuitive depiction of the frequency and prominence of specific terms within the dataset. The analysis underscores the pervasive connection of the majority of these data to everyday psychological counseling, further emphasizing the relevance and reasonability of the collected data.
@@
# Average Turns per Dialogue 5.90
# Average Tokens per Turn 136.28
# Average Tokens per Question 43.62
# Average Tokens per Answer 92.67
@@ III-B Multi-Turn Dialogue Construction
III-C knowledge-based QA Construction
In addition to the daily QA interactions, we also introduce some more abstract psychological knowledge data, such as explanations of psychological terms. We crawl books related to psychology from the web and then use Qwen-72B to extract knowledge-based QA from them. Specifically, we segment books into text spans using a predefined fixed length, identifying the nearest sentence or paragraph as segmentation indicators. These text spans serve as the fundamental units for subsequent QA generation through the utilization of LLMs. First, the LLM generates questions and their corresponding answers. These question-answer pairs are then input into two LLM-based student modules, one utilizing retrieval-augmented generation (RAG) and the other without RAG, to produce two new sets of answers. Subsequently, a teacher module, also based on an LLM, evaluates and selects the best answer from those generated by the student modules. Furthermore, to ensure the quality and accuracy of the generated QA pairs, a manual validation process is implemented, wherein human evaluators assess and eliminate low-quality data. In addition, we extract after-school exercises from several books along with their corresponding answer analyses. Finally, we obtain 10K knowledge-based QA data.